OpenStudy (lexi724):

fan and medal! A right triangle has a hypotenuse that measures 97 mm and one side that measure 72 mm. What is the length of the remaining side of the triangle?

OpenStudy (lexi724):

A.65 mm B.55 mm C.60 mm D.87 mm

OpenStudy (lexi724):

just need to know how to set it up correctly @iGreen @kidrah69

OpenStudy (igreen):

Use the Pythgorean theorem

OpenStudy (igreen):

\(a^2 + b^2 = c^2\) Where 'a' and 'b' are the two sides and 'c' is the hypotenuse. \(a^2 + 72^2 = 97^2\) Solve for 'a'

OpenStudy (lexi724):

__^2+5184=9409 9409-5184= 4225 So A=65^2 answer choice A

OpenStudy (bee_see):

correct